# A Survey on Sequential Recommendation

### v1

@@ 6.3.4 LLMs for Data Generation
LLMs can be leveraged to generate data for solving the data sparsity and cold-start problems. For example, KDALRD [159] leverages LLMs to describe relationships among items. Most existing models rely on predefined relations among items, which might result in sparse issues. To generate more item-to-item relations, KDALRD leverages LLMs to learn item representations. The item reconstruction module leverages the estimated relation and another item to reconstruct the target item. Finally, KDALRD leverages the predefined and generated relations for recommendation. FineRec [158] initially leverages LLMs to extract attribute pairs from reviews. Then by constructing an attribute-aware graph and aggregating, FineRec can obtain improved user and item embeddings for recommendation. For few-shot recommendation, ReLLa [160] can be used as a data augmentation method to augment data. ReLLa [160] leverages LLMs to retrieve the top- semantically relevant behaviors according to a target item. Leveraging the semantic knowledge of LLM, ReLLa enhances the quality of data. Finally, ReLLa is trained on a mixed dataset containing the original data and the retrieval-enhanced data.
@@ 6.4 Ultra-Long SR
An ultra-long behavior sequence typically refers to a scenario where the number of interacted items exceeds 1000. Compared with conventional behavior sequence modeling in SR models, long behavior sequence modeling presents several challenges. Firstly, modeling long sequences requires substantial computational resources. Secondly, long sequences often contain noise due to irrelevant behaviors. Finally, during the inference phase, modeling long sequences can be computationally expensive. For example, the time complexity of Transformer is . Therefore, SR models while modeling long sequences might face high latency. The general long sequence modeling methods are shown in Figure 13. Firstly, the most relevant items are retrieved from a whole interaction sequence. Then long-term interests are extracted from the retrieved items by an attention module. Short-term interests are extracted from recently interacted items by an attention module. The process of extracting the two interests is formalized as follows:
@@
Traditional SR methods focus on leveraging the most recent behaviors, while our UBR4CTR [163] emphasizes the most relevant behaviors. When only the recent interactions are leveraged, the long-term dependencies and periodicity might not be captured accurately. ETA-Net [164] aims to retrieve relevant items from long sequences in a more efficient manner. Instead of using a traditional dot-product operation, ETA-Net leverages a hashing-based target attention algorithm to calculate the Hamming distance, and then selects the top- relevant behavior items for a target item. Finally, ETA-Net leverages multi-head attention to learn long sequence embeddings for capturing diverse interests from the selected top- items. During prediction, ETA-Net leverages five different embeddings (i.e., long sequence embeddings, target item embeddings, short-sequence embeddings, user profile embeddings and context embeddings). For users with long interaction sequences, these sequences can be divided into sessions based on time intervals between interactions. As a result, HNUM [165] is designed with a two-layer architecture. The first layer is a session-level GRU model designed to capture users’ short-term preferences within the current session. The second layer is a user-level memory network, which can capture users’ long-term preferences across their entire interaction history. SIM [11] captures users’ interests using both general search unit (GSU) and exact search unit (ESU). GSU leverages both hard search and soft search methods. The hard search aims to retrieve interacted items belonging to the same category of a candidate item from a user’s long sequence. The soft search leverages inner product to identify top- relevant items. These top- items are used to construct users’ sub-sequences. By employing multi-head attention, ESU effectively models the relationship between candidate items and the sub-sequences to extract users’ long-term preferences. To solve the inconsistency between CP-GSU and ESU for target items, TWIN [166] leverages the same target-behavior relevance metric in both modules. Furthermore, to enhance efficiency, TWIN compresses user-item cross features into bias terms for target attention. To tackle ultra-long user behavior sequences, TWIN-V2 [167] leverages a hierarchical method to compress life-cycle behaviors. Similar items are classified into the same cluster. Meanwhile, TWIN-V2 leverages cluster-aware target attention to learn users’ diverse interests. R-RNN [168] learns a user’s global interests and recent interests based on global user behaviors and recent user behaviors, respectively. To improve efficiency while modeling long behavior sequences, SAM [169] leverages dual-query attention by regarding target item embedding and memory vector as queries. User interest center (UIC) is designed to decouple user interest modeling from the whole model. Modeling user interests needs to consume lots of computational time. UIC learns a user’s latest interest representations according to real-time user behavior events. MIMN [170] leverages memory utilization regularization and memory induction unit to learn multi-channel user interests. UBCS [171] proposes a behavior sampling module to sample short sequences based on candidate items, relevance and temporal information. To speed up the process of sampling, UBCS leverages an item clustering module to cluster candidate items for selecting some representative candidate items. CoFARS [172] takes into account the influence of contexts on user preferences. CoFARS leverages a prototype-based approach to identify contexts that reflect similar user preferences. Then, CoFARS constructs a temporal graph containing context nodes and prototype nodes for integrating temporal information into users’ interests. Finally, CoFARS identifies the prototypes that are consistent with the target context to generate a sub-sequence. Different from single-domain long sequential modeling, LCN [173] can transfer knowledge from a source domain to a target domain. In a cross representation production module, LCN leverages contrastive learning to bridge the connections among similar items in different domains. In a lifelong attention pyramid module, LCN leverages three kinds of cascading attentions to extract preferences from user lifelong behavior sequences considering candidate items. LRURec [174] leverages linear recurrent units and a recursive parallelization framework to enhance computational efficiency and realize low-cost inference. LinRec [175] reduces the quadratic time complexity of self-attention to a linear time complexity while preserving the advantages of attention mechanisms. To address the issue of sparse search behaviors failing to capture users’ interests, QIN [176] adopts a two-stage search method. Firstly, the top- items from a long sequence for user are retrieved using a pre-trained retrieval model. Secondly, QIN predicts whether user will purchase items based on the retrieved top- items. For QIN, it first searches sub-sequences based on the relevance between behaviors and queries. Secondly, QIN searches sub-sequences according to the relevance between behaviors and target items. By considering item ID field and attribute field separately, QIN leverages a fused attention unit to learn users’ representations better. To efficiently retrieve sub-sequences, SDIM [177] leverages an effective sampling method to generate sub-sequences. SDIM generates hash signatures for both the candidate item and each item in a user behavior sequence by multiple hash functions. SDIM can obtain the sub-sequences where items have the hash signatures the same as the candidate item. Item IDs might not exist in a dataset and different modal (i.e., text, images and attributes) embeddings are not aligned. Meanwhile, users’ interacted item sequence and users’ search query sequence reflect users’ interests from different perspectives. To address these problems, SEMINAR [178] leverages a pre-training search unit to learn query-item pairs. Then, SEMINAR leverages pre-trained embeddings to initialize the network. Additionally, SEMINAR leverages a multi-modal product quantization strategy to reduce time complexity. The basic framework of SEMINAR includes GSU and ESU. To model users’ long sequences and capture evolving users’ interests, HPMN [179] leverages memory slots to realize sequential patterns’ personalized memorization for each user. To capture multi-scale and evolving sequence patterns, HPMN leverages a hierarchical incremental updating mechanism.
@@ 6.5 Data-Augmented SR
Time intervals in users’ interaction sequences might vary significantly, which might fail to model users’ preferences due to preference drift. To solve this problem, TiCoSeRec [180] standardizes interaction sequences by transforming uneven time intervals into uniform distributions. TiCoSeRec consists of five operators (i.e., Ti-Crop, Ti-Reorder, Ti-Mask, Ti-Substitute and Ti-Insert). Different from traditional operators, the five data augmentation methods explicitly account for time intervals. Finally, TiCoSeRec leverages contrastive learning to ensure a high similarity between the augmented and original sequences. CL4SRec [12] leverages three kinds of data augmentation approaches (i.e., item crop, item mask and item reorder). Meanwhile, CL4SRec leverages contrastive learning between one augmented user interaction sequence and another augmented one. To alleviate the data sparsity problem and obtain stable users’ interests, TGCL4SR [181] leverages two methods (i.e., neighbor sampling and time disturbance) to obtain augmented subgraphs. The neighbor sampling constructs augmented subgraphs by randomly sampling the neighbor nodes of each item. The time disturbance constructs the hypergraph to randomly add some noise that obeys Gaussian distribution to timestamps. Finally, by using disturbed graph contrastive loss and subgraph contrastive loss, high-quality representations can be obtained. Driven by the potential of data-centric AI, DR4SR+ [182] incorporates a model-aware dataset personalizer to tailor the regenerated dataset. In the process of model-agnostic dataset regeneration, a regenerator is pre-trained by learning transition patterns obtained through rule-based methods in the original interaction sequences. DR4SR+ leverages a diversity-promoted regenerator to capture one-to-many mapping relationships between the original sequences and generated patterns. To generate personalized datasets for target items, DR4SR+ uses a dataset personalizer to evaluate the quality of each generated sample for a target model. Traditional data augmentation methods often modify an original interaction sequence randomly. For example, some items are masked in sequences randomly. These methods might break sequential correlations. To solve this problem, MStein [183] leverages Wasserstein discrepancy measurement to measure mutual information among augmented sequences and ensure the maximum of the mutual information. MBASR [184] considers users’ behaviors while data augmentation is conducted. DuoRec [185] leverages a contrastive learning regularizer to generate item embeddings with more uniform magnitude and frequency. END4Rec [186] improves the performance of Transformer by pruning noisy input. ASReP [187] leverages a pre-training Transformer to enhance the length of short users’ interaction sequences, and then the Transformer is fine-tuned by augmented sequences.
@@ 7 Empirical Studies
In this section, our survey will introduce some datasets, evaluation protocols and evaluation metrics used in sequential recommendation.
@@ 7.1 Datasets
Datasets | Multi-modal Features | Links |
MIND | text | https://msnews.github.io |
H-M | text, images | https://www.kaggle.com/c/h-and-m-personalized-fashion-recommendations/data |
Bili | text, images | https://github.com/westlake-repl/NineRec?tab=readme-ov-file |
Art of the Mix | text | https://brianmcfee.net/data/aotm2011.html |
LastFM | text | https://www.kaggle.com/datasets/harshal19t/lastfm-dataset |
Amazon 2023 | text, images, videos | https://huggingface.co/datasets/McAuley-Lab/Amazon-Reviews-2023 |
MicroLens | text, images, videos | https://github.com/westlake-repl/MicroLens |
MovieLens | text | https://grouplens.org/datasets/movielens/ |
Steam | text, images | https://huggingface.co/datasets/FronkonGames/steam-games-dataset |
Yelp | text, images | https://www.yelp.com/dataset |
Goodreads | text, images | https://www.kaggle.com/datasets/jealousleopard/goodreadsbooks |
Douban | text | https://www.kaggle.com/datasets/fengzhujoey/douban-datasetratingreviewside-information |
@@ 7.2 Evaluation Protocols
During the validation and testing phase, candidate items can be selected by three methods (i.e., random sampling, sampling based on popularity, and full sampling). Random sampling mainly selects a small number of candidate items randomly [1]. This method is simple yet effective. Because the SR models are usually deployed in ranking or re-ranking phase, there are only a small number of candidate items. The number of candidate items is consistent with the actual situation. However, random sampling might lead to variability in the candidate items. The evaluation results might be unstable with the varying candidate items. Sampling based on popularity is that the candidate items are selected based on popularity [103]. If the popularity of items is larger, the probability of these items that are selected as candidate items is larger. This method can reduce the influence of randomness. Users interact with popular items easily, therefore these items appear in candidate items more frequently. Compared with random sampling, this split method is more stable and provides a more accurate evaluation of recommendation performance. The final method considers all items in the dataset as candidate items [182]. This method can evaluate the recommendation performance of different models from a comprehensive view. However, it might consume lots of computational resources especially when the dataset contains a large number of items.

### v2

@@ 6.3.4 LLMs for Data Generation
LLMs can be leveraged to generate data for alleviating the data sparsity and cold-start problems. For example, KDALRD [166] leverages LLMs to describe relationships among items. Most existing models rely on predefined relations among items, which might result in sparse issues. To generate more item-to-item relations, KDALRD leverages LLMs to learn item representations. An item reconstruction module leverages the estimated relation and another item to reconstruct the target item. Finally, KDALRD leverages the predefined and generated relations for recommendation. FineRec [165] initially leverages LLMs to extract attribute pairs from reviews. Then by constructing an attribute-aware graph and aggregating, FineRec can obtain improved user and item embeddings for recommendation. For few-shot recommendation, ReLLa [167] can be used as a data augmentation method to augment data. ReLLa [167] leverages LLMs to retrieve the top- semantically relevant behaviors according to a target item. Leveraging the semantic knowledge of LLMs, ReLLa enhances the quality of data. Finally, ReLLa is trained on a mixed data containing the original data and the retrieval-enhanced data. LLM4DSR [152] leverages LLMs to identify noisy items and replace them with suggested ones.
@@ 6.4 Ultra-Long SR
An ultra-long behavior sequence typically refers to a scenario where the number of interacted items exceeds 1000. Compared with conventional behavior sequence modeling in SR models, ultra-long behavior sequence modeling presents several challenges. Firstly, modeling ultra-long sequences requires substantial computational resources. Secondly, ultra-long sequences often contain noise due to irrelevant behaviors. Finally, during the inference phase, modeling ultra-long sequences can be computationally expensive. For example, the time complexity of Transformer is . Therefore, SR models while modeling ultra-long sequences might face high latency. The general ultra-long sequence modeling methods are shown in Figure 13. Firstly, the most relevant items are retrieved from a whole interaction sequence. Then long-term interests are extracted from the retrieved items by an attention module. Short-term interests are extracted from recently interacted items by an attention module. The process of extracting the two interests is formalized as follows:
@@
Traditional SR methods focus on leveraging the most recent behaviors, while UBR4CTR [171] emphasizes the most relevant behaviors. When only the recent interactions are leveraged, the long-term dependencies and periodicity might not be captured accurately. ETA-Net [172] aims to retrieve relevant items from long sequences in a more efficient manner. Instead of using a traditional dot-product operation, ETA-Net leverages a hashing-based target attention algorithm to calculate the Hamming distance, and then selects the top- relevant behavior items for a target item. Finally, ETA-Net leverages multi-head attention to learn long sequence embeddings for capturing diverse interests from the selected top- items. During prediction, ETA-Net leverages five different embeddings (i.e., long sequence embeddings, target item embeddings, short-sequence embeddings, user profile embeddings and context embeddings). For users with long interaction sequences, these sequences can be divided into sessions based on time intervals between interactions. As a result, HNUM [173] is designed with a two-layer architecture. The first layer is a session-level GRU model designed to capture users’ short-term preferences within the current session. The second layer is a user-level memory network, which can capture users’ long-term preferences across their entire interaction history. SIM [11] captures users’ interests using both general search unit (GSU) and exact search unit (ESU). GSU leverages both hard search and soft search methods. The hard search aims to retrieve interacted items belonging to the same category of a candidate item from a user’s long sequence. The soft search leverages inner product to identify top- relevant items. These top- items are used to construct users’ sub-sequences. By employing multi-head attention, ESU effectively models the relationship between candidate items and the sub-sequences to extract users’ long-term preferences. To solve the inconsistency between CP-GSU and ESU for target items, TWIN [174] leverages the same target-behavior relevance metric in both modules. Furthermore, to enhance efficiency, TWIN compresses user-item cross features into bias terms for target attention. To tackle ultra-long user behavior sequences, TWIN-V2 [175] leverages a hierarchical method to compress life-cycle behaviors. Similar items are classified into the same cluster. Meanwhile, TWIN-V2 leverages cluster-aware target attention to learn users’ diverse interests. R-RNN [176] learns a user’s global interests and recent interests based on global user behaviors and recent user behaviors, respectively. To improve efficiency while modeling long behavior sequences, SAM [177] leverages dual-query attention by regarding target item embedding and memory vector as queries. User interest center (UIC) is designed to decouple user interest modeling from the whole model. Modeling user interests needs to consume lots of computational time. UIC learns a user’s latest interest representations according to real-time user behavior events. MIMN [178] leverages memory utilization regularization and memory induction unit to learn multi-channel user interests. UBCS [179] proposes a behavior sampling module to sample short sequences based on candidate items, relevance and temporal information. To speed up the process of sampling, UBCS leverages an item clustering module to cluster candidate items for selecting some representative candidate items. CoFARS [180] takes into account the influence of contexts on user preferences. CoFARS leverages a prototype-based approach to identify contexts that reflect similar user preferences. Then, CoFARS constructs a temporal graph containing context nodes and prototype nodes for integrating temporal information into users’ interests. Finally, CoFARS identifies the prototypes that are consistent with the target context to generate a sub-sequence. Different from single-domain long sequence modeling, LCN [181] can transfer knowledge from a source domain to a target domain. In a cross representation production module, LCN leverages contrastive learning to bridge the connections among similar items in different domains. In a lifelong attention pyramid module, LCN leverages three kinds of cascading attentions to extract preferences from user lifelong behavior sequences considering candidate items. LRURec [182] leverages linear recurrent units and a recursive parallelization framework to enhance computational efficiency and realize low-cost inference. LinRec [183] reduces the quadratic time complexity of self-attention to a linear time complexity while preserving the advantages of attention mechanisms. To address the issue of sparse search behaviors failing to capture users’ interests, QIN [184] adopts a two-stage search method. Firstly, the top- items from a long sequence for user are retrieved using a pre-trained retrieval model. Secondly, QIN predicts whether user will purchase items based on the retrieved top- items. For QIN, it first searches sub-sequences based on the relevance between behaviors and queries. Secondly, QIN searches sub-sequences according to the relevance between behaviors and target items. By considering item ID field and attribute field separately, QIN leverages a fused attention unit to learn users’ representations better. To efficiently retrieve sub-sequences, SDIM [185] leverages an effective sampling method to generate sub-sequences. SDIM generates hash signatures for both the candidate item and each item in a user behavior sequence by multiple hash functions. SDIM can obtain the sub-sequences where items have the hash signatures the same as the candidate item. Item IDs might not exist in a data and different modal (i.e., text, images and attributes) embeddings are not aligned. Meanwhile, users’ interacted item sequence and users’ search query sequence reflect users’ interests from different perspectives. To address these problems, SEMINAR [186] leverages a pre-training search unit to learn query-item pairs. Then, SEMINAR leverages pre-trained embeddings to initialize the network. Additionally, SEMINAR leverages a multi-modal product quantization strategy to reduce time complexity. The basic framework of SEMINAR includes GSU and ESU. To model users’ long sequences and capture evolving users’ interests, HPMN [187] leverages memory slots to realize sequential patterns’ personalized memorization for each user. To capture multi-scale and evolving sequence patterns, HPMN leverages a hierarchical incremental updating mechanism.
@@ 6.5 Data-Augmented SR
Time intervals in users’ interaction sequences might vary significantly, which might fail to model users’ preferences due to preference drift. To solve this problem, TiCoSeRec [188] standardizes interaction sequences by transforming uneven time intervals into uniform distributions. TiCoSeRec consists of five operators (i.e., Ti-Crop, Ti-Reorder, Ti-Mask, Ti-Substitute and Ti-Insert). Different from traditional operators, the five data augmentation methods explicitly account for time intervals. Finally, TiCoSeRec leverages contrastive learning to ensure a high similarity between the augmented and original sequences. CL4SRec [12] leverages three kinds of data augmentation approaches (i.e., item crop, item mask and item reorder). Meanwhile, CL4SRec leverages contrastive learning between one augmented user interaction sequence and another augmented one. To alleviate the data sparsity problem and obtain stable users’ interests, TGCL4SR [189] leverages two methods (i.e., neighbor sampling and time disturbance) to obtain augmented subgraphs. The neighbor sampling constructs augmented subgraphs by randomly sampling the neighbor nodes of each item. The time disturbance constructs the hypergraph to randomly add some noise that obeys Gaussian distribution to timestamps. Finally, by using disturbed graph contrastive loss and subgraph contrastive loss, high-quality representations can be obtained. Driven by the potential of data-centric AI, DR4SR+ [190] incorporates a model-aware data personalizer to tailor the regenerated data. In the process of model-agnostic data regeneration, a regenerator is pre-trained by learning transition patterns obtained through rule-based methods in the original interaction sequences. DR4SR+ leverages a diversity-promoted regenerator to capture one-to-many mapping relationships between the original sequences and generated patterns. To generate a personalized data for target items, DR4SR+ uses a data personalizer to evaluate the quality of each generated sample for a target model. Traditional data augmentation methods often modify an original interaction sequence randomly. For example, some items are masked in sequences randomly. These methods might break sequential correlations. To solve this problem, MStein [191] leverages Wasserstein discrepancy measurement to measure mutual information among augmented sequences and ensures the maximum of the mutual information. MBASR [192] considers users’ behaviors while data augmentation is conducted. DuoRec [193] leverages a contrastive learning regularizer to generate item embeddings with more uniform magnitude and frequency. END4Rec [194] improves the performance of Transformer by pruning noisy input. ASReP [195] leverages a pre-training Transformer to enhance the length of short users’ interaction sequences, and then the Transformer is fine-tuned by augmented sequences.
@@ 7 Empirical Studies
In this section, our survey will introduce some datasets, evaluation protocols and evaluation metrics used in sequential recommendation. Finally, we will present experimental results of some representative SR models.
@@ 7.1 Datasets
| Datasets | Multi-modal Features |
| MIND888https://msnews.github.io | text |
| H-M999https://www.kaggle.com/c/h-and-m-personalized-fashion-recommendations/data | text, images |
| Bili101010https://github.com/westlake-repl/NineRec?tab=readme-ov-file | text, images |
| Art of the Mix111111https://brianmcfee.net/data/aotm2011.html | text |
| LastFM121212https://www.kaggle.com/datasets/harshal19t/lastfm-dataset | text |
| Amazon 2023131313https://huggingface.co/datasets/McAuley-Lab/Amazon-Reviews-2023 | text, images, videos |
| MicroLens141414https://github.com/westlake-repl/MicroLens | text, images, videos |
| MovieLens151515https://grouplens.org/datasets/movielens/ | text |
| Steam161616https://huggingface.co/datasets/FronkonGames/steam-games-dataset | text, images |
| Yelp171717https://www.yelp.com/dataset | text, images |
| Goodreads181818https://www.kaggle.com/datasets/jealousleopard/goodreadsbooks | text, images |
| Douban191919https://www.kaggle.com/datasets/fengzhujoey/douban-datasetratingreviewside-information | text |
@@ 7.2 Evaluation Protocols
During the validation and testing phase, candidate items can be selected by three methods (i.e., random sampling, sampling based on popularity, and full sampling). Random sampling selects a small number of candidate items randomly [1]. This method is simple yet effective. Because the SR models are usually deployed in ranking or re-ranking phase, there are only a small number of candidate items. The number of candidate items is consistent with the actual situation. However, random sampling might lead to variability in the candidate items. The evaluation results might be unstable with the varying candidate items. Sampling based on popularity is that the candidate items are selected based on popularity [103]. If the popularity of items is larger, the probability of these items that are selected as candidate items is larger. This method can reduce the influence of randomness. Users interact with popular items easily, therefore these items appear in candidate items more frequently. Compared with random sampling, this split method is more stable and provides a more accurate evaluation of recommendation performance. The final method considers all items in the dataset as candidate items [190]. This method can evaluate the recommendation performance of different models from a comprehensive view. However, it might consume lots of computational resources especially when the dataset contains a large number of items.
@@ 7.3 Evaluation Metrics
7.4 Experimental Results
| Models | Office | Game | Toy | |||
| HR@10 | NDCG@10 | HR@10 | NDCG@10 | HR@10 | NDCG@10 | |
| SASRec | 0.1056 | 0.0710 | 0.0186 | 0.0093 | 0.0561 | 0.0312 |
| BERT4Rec | 0.0825 | 0.0634 | - | - | 0.0352 | 0.0179 |
| FDSR | 0.1118 | 0.0868 | 0.0190 | 0.0101 | 0.0568 | 0.0344 |
| S3-Rec | 0.1030 | 0.0653 | 0.0195 | 0.0094 | 0.0538 | 0.0276 |
| UniSRec w/o IDs | 0.1013 | 0.0619 | - | - | - | - |
| UniSRec | 0.1280 | 0.0831 | 0.0225 | 0.0115 | - | - |
| MISSRec w/o IDs | 0.1258 | 0.0795 | - | - | - | - |
| MISSRec | 0.1301 | 0.0842 | - | - | - | - |
| RECFORMER | - | - | 0.0243 | 0.0114 | - | - |
| TIGER | - | - | 0.0222 | 0.0114 | - | - |
| SpecGR | - | - | 0.0229 | 0.0115 | - | - |
| P5 | - | - | - | - | 0.0543 | 0.0356 |
| POD | - | - | - | - | 0.0566 | 0.0421 |
| Lite-LLM4Rec | - | - | - | - | 0.0682 | 0.0488 |
To show the relative recommendation performance of different SR models, we quote and copy the results of some recent works [138, 147, 149]. The compared SR models can be classified into five categories, including pure ID-based SR (i.e., SASRec [1] and BERT4Rec [36]), SR with side information (i.e., FDSA [121] and S3Rec [120]), multi-modal SR (i.e., UniSRec [8], MISSRec [138] and RECFORMER [137]), generative SR (i.e., TIGER [141] and SpecGR [147]), and LLM-powered SR (i.e., P5 [148], POD [150] and Lite-LLM4Rec [149]). “Office”, “Game” and “Toy” are three domains in an Amazon dataset. Two widely-used metrics (i.e., HR@10 and NDCG@10) are adopted for performance evaluation. In each domain, the leave-one-out strategy is leveraged to obtain the training, validation and test data. Meanwhile, all items are taken as the candidate ones in recommendation. The data in each domain is processed according to the corresponding papers [138, 147, 149]. From the experimental results in Table 5, we can have the following observations:
SR models with side information achieve better performance than pure ID-based SR ones, which demonstrates that leveraging item features can alleviate the cold-start and sparsity problem to some extent.
By removing item IDs from multi-modal SR models, the recommendation performance will decrease. It is because that the item IDs can capture the collaborative signals and users’ fine-grained preferences.
LLM-powered SR models show better recommendation performance than traditional SR models. We can see that the semantic knowledge extracted by LLMs are useful for enhancing the recommendation performance.
Compared with pure ID-based and multi-modal SR models, generative SR models show comparable and even better performance. By leveraging semantic IDs, generative models can capture different grained semantic information.